OpenStudy (anonymous):

Whats 16 in 75 as a fraction? Its either 1/20, 1/5, 1/8, 2/3.

OpenStudy (anonymous):

it is none of those \(75=3\times 5^2\) and \(16=2^4\) there are no common factors, nothing to reduce